OpenStudy (anonymous):

Naomi is moving two boxes (A = 20kg and B = 10kg), one on top of the other, with magnitude F = 150 N along the horizontal. The coefficient of kinetic friction between box A and the floor is 0.40 while the coefficient of static friction between box A and box B is 0.80. What is the acceleration of the boxes. a. 1.1 m/s2 c. 5.0 m/s2 b. 8.9 m/s2 d. 2.9 m/s2

OpenStudy (anonymous):

please help me @ hoslos

OpenStudy (anonymous):

@helpme25 @hoslos

OpenStudy (anonymous):

First of all, to get the exact acceleration n of the boxes, the force applied has to be the one that will not make them slide over each other. To do that, we need to work with the upper box(B): Normal contact force or N= mg or 10*9.81=98.1N Ff(frictional force) = required applied force to avoid sliding=€N or 0.8*98.1=78.48N Do you have the diagram?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

i have sorry

OpenStudy (anonymous):

Hehe, ok. Please post it out.

OpenStudy (anonymous):

Do they mention anything about sliding or no sliding of boxes?

OpenStudy (anonymous):

nope :D

OpenStudy (anonymous):

Guess not. ALRIGHt forget about what I wrote ealier. Let us go for the case of the boxes and floor. First we find the frictional force, find N = mg or (10+20)*9.81=294.3N The frictional force is then $N or 0.4*294.3=117.72N. Use the 2: law of motion's formulaof F-R=ma or 150-117.72=(20+10)a and a is 1.076, approximately 1.1m/s^2

OpenStudy (anonymous):

where did u get the R?

OpenStudy (anonymous):

R is the resistive force or frictional force that I calculated right there.